### v1

# On the Dimension of Pullback Attractors in Recurrent Neural Networks

Abstract
Recurrent Neural Networks (RNNs) are high-dimensional state space models capable of learning functions on sequence data. Recently, it has been conjectured that reservoir computers, a particular class of RNNs, trained on observations of a dynamical systems can be interpreted as embeddings. This result has been established for the case of linear reservoir systems. In this work, we use a nonautonomous dynamical systems approach to establish an upper bound for the fractal dimension of the subset of reservoir state space approximated during training and prediction phase. We prove that when the input sequences comes from an $N_{\text{in}}$-dimensional invertible dynamical system, the fractal dimension of this set is bounded above by $N_{\text{in}}$. The result obtained here are useful in dimensionality reduction of computation in RNNs as well as estimating fractal dimensions of dynamical systems from limited observations of their time series. It is also a step towards understanding embedding properties of reservoir computers.
1 Introduction
Reservoir computing (RC) framework has offered a computationally efficient way to train recurrent neural networks by randomly generating the internal weights and keeping them fixed while focusing training only on the output layer [18, 25]. This allows RC systems to avoid the famous vanishing and exploding gradient problem [26], and are one of the main contenders for biologically plausible temporal learning [8, 9]. Furthermore, large class of RC systems have been shown to be universal approximators of fading memory mappings on left infinite input sequences [14, 11]. Consequently, even with the restriction on training, RC systems achieve state-of-the-art performance on a number of benchmark including as time-series classification [16], systems identification [19], adaptive filtering [34], attractor reconstruction [24] etc.
Beyond the theory of universal approximations, it is unclear how RC systems are able to achieve such performance on various benchmark tasks. In past few years, there has been research to understand how RC systems process inputs and make their decisions. In the case of sequence-to-sequence classifications, it has been shown for example in [31, 3] that reservoir computers trained with output feedback take advantage of stable fixed points and their relative locations to solve such tasks. As well as in the continuous output case, errors in RC systems trained on such tasks have been identified with the emergence of spurious/untrained attractors [10, 3]. Transitions to such attractors may cause errors over short time steps in sequence-to-sequence classification problems and over longer time steps in continuous output problems [10].
In [15], the authors conjectured that reservoir computers trained on generic continuous observations of a smooth invertible dynamical system on an $m$-dimensional compact manifold embeds the dynamics of the dynamical system (not the observations) it is state space. This has been settled in the case of linear reservoirs [13]. If this conjecture holds in the general case, it will have significant impact on how we understand computation in RC systems. First, the embedding of the dynamics on the manifold on to reservoir space is achieved by a unique map that sends points on the manifold to a corresponding point on the pullback attractor – a nonautonomous attractive invariant set in the reservoir space. As universal approximation results are established with respect to the pullback attractor in the reservoir space, this result means that an RC system provides a faithful higher dimensional representation of the dynamical system and allows learning to be done more easily. Secondly, As there is a reservoir computer whose dynamics corresponds to the $(2m+1)$-time delay map, this result would be a generalisation of Takens’ delay embedding theorem [32] or the prevalent versoin of Sauer [30]. Consequently, this would mean reservoir systems should preserve dynamic invariance such as Lyapunov exponents and fractal dimensions.
There are two main ways to proving an embedding theorem for dynamical systems – there is the Takens’ approach [32] or the prevalence version of Sauer et al. [30]. In the both cases, the question of dimensions and embeddings are inextricably linked. In this work, we consider RC systems in a broad class of input-driven systems of the form
$x_{n}=g(u_{n},x_{n-1})$ (1)
where $g$ is some function on $U$ (the space of inputs) and $X$ (the reservoir state space). We investigate the fractal dimension input-driven systems of the type (1) using nonautonomous dynamical systems theory with particular focus on more general space of input signals than just finite dimensional dynamical systems. We show that under mild conditions on $g$, the fractal dimension of the pullback attractor associated with such systems is bounded above by the fractal dimension of the state of the input sequences $\mathcal{U}=U^{\mathbb{Z}}$ in the product topology. In the case that the set $U$ contains an open set, we show that pullback attractor also contain an open set in the reservoir space and thus has the same dimension as that of the driven reservoir. As a special case, we consider input sequences coming from an invertible $N_{\text{in}}$-dimensional dynamical system $\mathcal{V}$. Considering $\mathcal{V}$ as a subset of $({\mathbb{R}^{N_{\text{in}}}})^{\mathbb{Z}}$ in the product, we show that fractal dimension of $\mathcal{V}$ in this topology is bounded above by $N_{\text{in}}$.
The remainder of this paper is organised as follows. Section 2 describes notations of this paper. We then introduce RNNs as nonautonomous dynamical system and provide preliminary result from the theory of nonautonomous dynamical systems. Section 3 contains the main contribution of this paper. We first showed in section 3.1, following results from [15], that the pullback attractor of a state contracting RC system is the image of a continuous map with respect to the product topology on the space of input sequences. We provide alternative interpretations of this map as well as those offered in the literature. In section 3.2, we outline existing results and conditions on the theory of fractal dimension of pullback attractors [5]. We then use the map obtained in 3.1 to establish a bound for the fractal dimension of pullback attractors.
2 Background
Suppose $N_{\text{in}}\in\mathbb{N}$ and $U$ is compact and the closure of an open subset of $\mathbb{R}^{N_{\text{in}}}$ where $\mathbb{R}^{N_{\text{in}}}$ is equipped with the standard Euclidean metric
$d(u,v)=\sum_{i=1}^{N_{\text{in}}}(u_{i}-v_{i})^{2}\quad u,v\in\mathbb{R}^{\mathbb{N_{\text{in}}}}.$
We will take $V$ to be the interior $\text{int}\>U$ of $U$ $i.e$ the largest open set contained in $U$. We consider the Cartesian product $\mathcal{U}=U^{\mathbb{Z}}$ consisting of all bi-infinite sequences of $N_{\text{in}}$-dimensional vectors $\mathbf{u}=\{\mathbf{u}[n]\}_{n=-\infty}^{\infty}$ where the terms $\mathbf{u}[n]\in U$ is the $n^{\text{th}}$ of $\mathbf{u}$ for each $n\in\mathbb{Z}$. $\mathcal{U}$ is considered as topological space equipped with the product topology induced by the metric
$D(\mathbf{u},\mathbf{v})=\sup_{n\in\mathbb{Z}}\{2^{-|n|}d(\mathbf{u}[n],\mathbf{v}[n])\}.$
Roughly speaking two input sequences are close in this topology if they are close at only finite number of terms and maybe arbitrarily far away from each other at other terms. By Tychonoff’s theorem [21, Theorem 37.3], $\mathcal{U}$ is compact in this topology. Another topology on $\mathcal{U}$ is the uniform topology induced by the metric
$D^{\text{unif}}(\mathbf{u},\mathbf{v})=\sup_{n\in\mathbb{Z}}\{d(\mathbf{u}[n],\mathbf{v}[n])\}.$
Unless otherwise stated, we will consider the product topology on $\mathcal{U}$.
We denote the set $U^{\mathbb{Z}_{-}}$ consisting of all left-infinite input sequences taking values in $U$ as $\mathcal{U}^{-}$ and the set $U^{\mathbb{Z}_{+}}$ of all right infinite input sequences taking values in $U$ as $\mathcal{U}^{+}$. We denote $C(\mathcal{U},X)$ as the space of continuous functions from $\mathcal{U}$ (endowed with the product topology) to $X$. Equipped with the uniform metric
$\rho(H,F)=\sup_{\mathbf{u}\in\mathcal{U}}d(H(\mathbf{u}),F(\mathbf{u}))\quad H,F\in C(\mathcal{U},X),$ (2)
$C(\mathcal{U},X)$ is a complete metric space [21, Theorem 43.6].
Given an initial condition $\mathbf{x}[0]\in\mathbb{R}^{N_{r}}$ with $N_{r}\in\mathbb{N}$, the dynamics of the internal states of a RNN with $N_{r}$ hidden units driven by an input sequence $\mathbf{u}\in\mathcal{U}^{+}$ is given by
$\mathbf{x}[n+1]$ $=g{(W\mathbf{x}[n]+W^{\text{in}}\mathbf{u}[n+1]+W^{\text{fb}}\mathbf{y}[n])}$ (3)
$\mathbf{y}[n+1]$ $=h(W^{\text{out}}\mathbf{x}[n+1]),\quad n\in\mathbb{N}_{0}$ (4)
where $W\in\mathbb{R}^{N_{r}\times N_{r}}$ is called the recurrent matrix, $W^{\text{in}}\in\mathbb{R}^{N_{r}\times N_{\text{in}}}$ is the input-to-network coupling, $W^{\text{fb}}\in\mathbb{R}^{N_{r}\times N_{\text{out}}}$ feeds back the output into the network and $W^{\text{out}}\in\mathbb{R}^{N_{r}\times N_{\text{out}}}$ is the readout weights. The functions $h$ and $g$ are activation functions applied component-wise. The activation functions $g$ is usually taken to be the hyperbolic tangent $\tanh$ and $h$ the identity. In the case that $h$ is the identity, the dynamics reduces to
$\mathbf{x}[n+1]$ $=g{(\tilde{W}\mathbf{x}[n]+W^{\text{in}}\mathbf{u}[n+1])}$ (5)
$\mathbf{y}[n+1]$ $=W^{\text{out}}\mathbf{x}[n+1]$ (6)
where $\tilde{W}=W+W^{\text{fb}}W^{\text{out}}$ is the effective recurrent matrix. RNNs can also be adapted for multiple time-scales [33] using the Leaky integrator RNNs with dynamics
$\mathbf{x}[n+1]$ $=(1-\alpha)\mathbf{x}[n]+\alpha g{(W\mathbf{x}[n]+W^{\text{in}}\mathbf{u}[n+1]+W^{\text{fb}}\mathbf{y}[n]}$ (7)
at leak rate $\alpha$ [20]. In all three cases, the current state $\mathbf{x}[n+1]$ of an RNN depends on the previous state $\mathbf{x}[n]$ and the current input $\mathbf{u}[n+1]$. This observation will be important when we consider RNNs as nonautonomous dynamical systems.
2.1 RNNs as nonautonomous dynamical system
The input-driven dynamics of RNNs is best understood in the lens of nonautonomous dynamical systems [22]. There are two main formulations for dealing with nonautonomous dynamical systems: process formulation and the skew product flow formulation or the cocycle formulation [^1] Although the system of processes formulation of Vishik and Chepyzhov [5] is sometimes easier to work with (See [7]).. The process formulation is useful when dealing with the properties of a single input sequence while the skew product flow formulation allows more general treatment of all input sequence taking values in a compact set. Following [4], we will use the skew product flow formulation in this paper.
We define the canonical projection
$\pi_{n}:(\mathbb{R}^{N_{\text{in}}})^{\mathbb{Z}}$ $\to\mathbb{R}^{N_{\text{in}}},$
$\pi_{n}(\mathbf{u})$ $=\mathbf{u}[n].$
Let
$\theta:\mathcal{U}$ $\to\mathcal{U}$
$\{\mathbf{u}[n]\}_{n=-\infty}^{\infty}$$\to\{\mathbf{u}[n+1]\}_{n=-\infty}^{\infty}$
be the left shift operator. The left shift operator is continuous (with respect to the product topology) since each of its projections $\pi_{n}(\theta(\mathbf{u}))=\mathbf{u}_{n+1}$ is continuous as a function from $\mathcal{U}$ to $U$. Furthermore
$D(\theta(\mathbf{u}),\theta(\mathbf{v}))$ $=\sup_{n\in\mathbb{Z}}2^{-n}d(\pi_{n}(\theta(\mathbf{u})),\pi_{n}(\theta(\mathbf{v})))$
$=\sup_{n\in\mathbb{Z}}2^{-n}d(\mathbf{u}[n+1],\mathbf{v}[n+1])$
$=2\sup_{n\in\mathbb{Z}}2^{-(n+1)}d(\mathbf{u}[n+1],\mathbf{v}[n+1])$
$=2D(\mathbf{u},\mathbf{v}).$
It is also invertible with inverse $\theta^{-1}$ as the right shift operator satisfying $\pi_{n}(\theta^{-1}(u))=\mathbf{u}[n-1]$ which is also continuous. It can easily be checked that $\theta\circ\theta^{-1}$ is the identity on $\mathcal{U}$ and that
$D(\theta^{-1}(\mathbf{u}),\theta^{-1}(\mathbf{v}))=\frac{1}{2}D(\mathbf{u},\mathbf{v}).$
Thus $(\theta,\mathcal{U})$ is a discrete time dynamical system with $\theta^{0}(\mathbf{u})=\mathbf{u}$ and
$\theta^{m+n}(\mathbf{u})=\theta^{m}\circ\theta^{n}(\mathbf{u})\quad\text{for all }m,n\in\mathbb{Z}\>\text{and for all }\>\mathbf{u}\in\mathcal{U}.$
Definition 2.1 (Skew product flow, [22]).
Suppose $X\subseteq\mathbb{R}^{N_{r}}$ is a compact and the closure of an open subset of $\mathbb{R}^{N_{r}}$. Let $g:U\times X\to\mathbb{R}^{N_{r}}$ be a continuous map. A cocycle mapping $\Phi:\mathbb{N}_{0}\times U\times\mathbb{R}^{N_{r}}\to\mathbb{R}^{N_{r}}$ is defined as
$\Phi(0,\mathbf{u},x):=x\quad\forall\mathbf{u}\in\mathcal{U}\>\text{ and }\>\forall x\in\mathbb{R}^{N_{r}},$ (8)
$\Phi(n,\mathbf{u},x):=g(\pi_{0}(\theta_{n}(\mathbf{u})),\Phi(n-1,\mathbf{u},x))=g(\mathbf{u}[n],\Phi(n-1,\mathbf{u},x))\quad\forall n\in\mathbb{N}.$ (9)
The pair $(\theta,\Phi)$ is called a skew product flow with base space $\mathcal{U}$ and state space $X$.
Remark 1.
The assumption of compactness on the state space $X$ is a natural one at least when dealing with recurrent neural networks with squashing activation functions i.e their image is contained in a compact set. Even in the case of leaky-integrator RNNs, it can be shown that their is compact invariant set that eventually contains the entire forward dynamics [4, Proposition B.1].
The map $\Phi$ is called a cocycle because it satisfies the cocycle property. The proof of this is relatively straightforward and is given in the appendix for completeness.
Lemma 2.1.
The map $\Phi$ satisfies the cocycle property: For any $\mathbf{u}$ and $x_{0}\in X$,
$\Phi(n+m,\mathbf{u},x_{0})=\Phi(n,\theta^{m}(\mathbf{u}),\Phi(m,\mathbf{u},x_{0}))\quad\text{for all }m,n\in\mathbb{N}_{0}.$ (10)
2.1.1 Attractors, Entire Solutions and the Echo State Property
As explained in [22, Chapter 3], the behaviour of nonautonomous dynamical systems is best understood using an indexed family $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ of subsets of $X$. For each $\mathbf{u}\in\mathcal{U}$, the set $\mathbf{A}_{\mathbf{u}}$ is called the $\mathbf{u}$-fiber of $\mathbf{A}$ and the set $\mathbf{A}$ is called a nonautonomous set. $\mathbf{A}$ is said to be non-empty (resp. compact) if each of the fibers are non-empty (resp. compact). Given a non-empty and compact nonautonomous set $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\{\mathbf{u}\in\mathcal{U}\}}$, we say that $\mathbf{A}$ is $\Phi$-positively invariant if for any $\mathbf{u}\in\mathcal{U}$
$\Phi(n,\mathbf{u},\mathbf{A}_{\mathbf{u}})\subseteq\mathbf{A}_{\theta^{n}(\mathbf{u})}\quad\text{for all }\>n\in\mathbb{N},\mathbf{u}\in\mathcal{U}.$ (11)
Analogously, we say that $\mathbf{A}$ is $\Phi$-invariant if
$\Phi(n,\mathbf{u},\mathbf{A}_{\mathbf{u}})=\mathbf{A}_{\theta^{n}(\mathbf{u})}\quad\text{for all }\>n\in\mathbb{N},\mathbf{u}\in\mathcal{U}.$ (12)
A sequence $\{\mathbf{x}[n]\}_{n\in\mathbb{Z}}$ is called an entire solution for $\mathbf{u}$ if
$\mathbf{x}[n]=\Phi(n-m,\theta^{m}(\mathbf{u}),\mathbf{x}[m])\quad\text{ for all }m,n\in\mathbb{Z},n\geq m.$
If a non-empty and compact nonautonomous set is $\Phi$-positively invariant, then there exist at least one entire solutions associated to every input sequence [22, Chapter 2, Lemma 2.20]. Note that, an entire solution, if it exist need not be unique. This generally is the case if for a given fixed $u\in U$, $g(x,u)$ has more than one fixed point. For example if for some $u\in U$ fixed, $z_{0},x_{0}\in X$ are fixed points of the map $g(x,u)$ with $x_{0}\neq z_{0}$, then there are infinitely many entire solution $\mathbf{x}$ one can assign to the input sequence $\mathbf{u}=\{u\}$ with $\mathbf{x}[0]=x_{0}$. In particular, while forward uniqueness is guaranteed for any well-define $g$, we may not always have backward uniqueness.
Although there are various generalisation of attractors to the nonautonomous settings, most of the results from autonomous dynamical systems have similar counterparts when the pullback definition (given below) is considered. In the what follows,
$\text{dist}(A,B):=\sup_{a\in A}{\inf_{b\in B}d(a,b)}$
is the Hausdorff semi-metric between two non-empty subsets $A$, $B$ of a metric space $(X,d)$. It can be easily checked that if $\text{dist}(A,B)=0$, then $\overline{A}\subseteq\overline{B}$ (See [28, Chapter 1]).
Definition 2.2 (Pullback Attractor).
A nonautonomous set $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ is called the pullback attractor of the skew-product flow $(\theta,\Phi)$ if for each input $\mathbf{u}\in\mathcal{U}$ and for each $n\in\mathbb{Z}$ and for every bounded subset $B$ of $X$
$\lim_{n\to\infty}\text{dist}(\Phi(n,\theta^{-n}(\mathbf{u}),B),\mathbf{A}_{\mathbf{u}})=0.$
We say that $\mathbf{A}$ is uniformly pullback attracting if
$\lim_{n\to\infty}\sup_{\mathbf{u}\in\mathcal{U}}\text{dist}(\Phi(n,\theta^{-n}(\mathbf{u}),B),\mathbf{A}_{\mathbf{u}})=0.$
An alternative notion of attraction in the nonautonomous setting is forward attraction.
Definition 2.3.
A nonautonomous set $\mathbf{A}=\{\mathbf{F}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ is called a forward attractor for the skew product flow $(\theta,\Phi)$ if for any input $\mathbf{u}\in\mathcal{U}$ and any bounded subset $B$ of $\mathbb{R}^{N_{r}}$, we have that
$\lim_{n\to\infty}\text{dist}(\Phi(n,\mathbf{u},B),\mathbf{F}_{\theta^{n}(\mathbf{u})})=0.$
$\mathbf{F}$ is called a uniform forward attractor if
$\lim_{n\to\infty}\sup_{\mathbf{u}\in\mathcal{U}}\text{dist}(\Phi(n,\mathbf{u},B),\mathbf{F}_{\theta^{n}(\mathbf{u})})=0$
for any bounded subset $B$ of $\mathbb{R}^{N_{r}}$.
Remark 2.
A Pullback attractor is in generally not a forward attractor. However, any uniform pullback attractor is also a uniform forward attractor and vice versa. To see this, fix $n\in\mathbb{Z}$ and note that
$\sup_{\mathbf{u}\in\mathcal{U}}\text{dist}(n,\theta^{-n}(\mathbf{u}),B),\mathbf{A}_{n})=\sup_{\mathbf{u}\in\mathcal{U}}\text{dist}(2n,\mathbf{u},B),\mathbf{A}_{\theta^{n}(\mathbf{u})})$
and then take $n\to\infty$.
Remark 3.
It is important to remark that there is another definition of pullback and forward attraction suitable for one input sequence. These definitions are essentially equivalent to the pullback/forward attraction with respect to the process formulation. Since we are concern with all input sequences in $\mathcal{U}$, we use the definition of pullback and forward attraction suitable for this case (See [22, Chapter 3] for more details).
As we already made the assumption that the state space $X$ is compact, the fibers of a pullback attractor are contained in a bounded set. From this, one can conclude that the pullback attractor is unique.
Theorem 2.2.
Suppose that $\mathbf{A}=\{\mathbf{A}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ and $\mathbf{F}=\{\mathbf{F}_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ are pullback attractor for $(\theta,\Phi)$. Then $\mathbf{F}_{\mathbf{u}}=\mathbf{A}_{\mathbf{u}}$.
Proof.
Since $X$ is compact, the set $Y=\overline{\bigcup_{\mathbf{v}\in\mathcal{U}}\mathbf{F}_{\mathbf{v}}}\subseteq X$ is bounded. Thus
$\text{dist}(\mathbf{F}_{\mathbf{u}},\mathbf{A}_{\mathbf{u}})\leq\lim_{n\to\infty}\text{dist}(\Phi(n,\theta^{-n}(\mathbf{u}),Y),\mathbf{A}_{\mathbf{u}})=0$
from which it follows that $\mathbf{F}_{\mathbf{u}}\subseteq\mathbf{A}_{\mathbf{u}}$. Similarly $\mathbf{A}_{\mathbf{u}}\subseteq\mathbf{F}_{\mathbf{u}}$. ∎
Another important consequence of compactness on $X$ and $U$ is that we have the following characterisation [2, Corollary 1.18] of the pullback attractor $\{\mathbf{A}_{\mathbf{u}}\}$ as
$\mathbf{A}_{\mathbf{u}}=\{x\in X:\exists\>\text{an entire solution }\mathbf{x}\text{ associated to }\mathbf{u},n\in\mathbb{Z}\text{ satisfying }\>\mathbf{x}[n]=x\}.$
As pointed out earlier, existence of a unique forward trajectory does not guarantee a unique backward trajectory for a given input sequence. The property that every input sequence has a unique entire solution associated to it is called the echo state property.
Definition 2.4 (Echo state property).
The skew product flow $(\theta,\Phi)$ is said to have the echo state property if for any input sequence $\mathbf{u}$, there exist a unique entire solution associated to it.
If a cocycle satisfies the echo state property, the $\mathbf{u}$-fiber of the pullback attractors is exactly the entire solution associated to $\mathbf{u}$ since pullback attractors are composed of entire solutions.
3 Theoretical Results
In this section, we will show that under certain conditions, the fibers of the pullback attractor can be obtained as the image of a continuous map $H$ from $\mathcal{U}$ to $\mathbb{R}^{N_{r}}$. This result has appeared for example in [15, 12] as the echo state mapping theorem in the context of manifolds. It is known that this map is also continuously differentiable if one considers the base space as a manifold and if the dynamics on this manifold is also continuously differentiable. However, we would not seek such regularity on the dynamics of input sequences. In section 3.2, this map will be useful in showing that the box-counting dimension of the subset of $\mathbb{R}^{N_{r}}$ that contains the pullback attractor associated with input sequences $\mathbf{u}$ contained in $\mathcal{V}\subseteq\mathcal{U}$ is bounded above by the box-counting dimension of $\mathcal{V}$ in the product topology, and the box-counting dimension of $H(\mathcal{U})$ is always $N_{r}$.
3.1 Pullback Attractor as the image of a continuous map
Let $C(\mathcal{U},X)$ be the space of continuous functions from $\mathcal{U}$ to $X$ endowed with the uniform metric (2). Let $H_{0}\in C(\mathcal{U},X)$ and define
$H_{n}(\mathbf{u})=\Phi(1,\mathbf{u},H_{n-1}(\theta^{-1}(\mathbf{u})))\quad\text{for }\>n\in\mathbb{N}$
The sequence of maps $\{H_{0},H_{1},\dots,H_{n}\}$ is called the generalised echo state family of length $n$, following the notation of [15]. The goal of this section is to generalise the result of [15] that if $g:U\times X\to X$ is uniformly contracting in the state variable i.e there exist $\mu<1$ such that for all $x,z\in X$ and for all $u\in U$,
$||g(u,x)-g(u,z)||\leq\mu||x-z||,$
then $H_{n}$ converge to a unique continuous map from $\mathcal{U}$ to $X$ no matter the choice of initial $H_{0}$. The map $H$ also naturally relates to the pullback attractors of the skew product $(\theta,\Phi)$.
Theorem 3.1.
Suppose that the map $g:U\times X\to X$ is uniformly a $\mu$-contraction in the state variable. Then there exists a unique continuous map $H\in C(\mathcal{U},X)$ such that $\lim_{n\to\infty}H_{n}=H$ with respect to the uniform metric $\rho$ on $C(\mathcal{U},X)$ for all initial choice $H_{0}\in C(\mathcal{U},X)$.
Proof.
The proof will follow arguments of the proof of [15, Theorem 2.2.2]. Let $\mathbf{\Psi}:C(\mathcal{U},X)\to C(\mathcal{U},X)$ be the map $H\to\Phi(1,\theta^{0}(\cdot),H(\theta^{-1}(\cdot))$. We will show that $\mathbf{\Psi}$ is a $\mu$-contraction on $C(\mathcal{U},X)$. Let $F,H\in C(\mathcal{U},X)$. Then
$\rho(\mathbf{\Psi}(F),\mathbf{\Psi}(H))$ $=\sup_{\mathbf{u}\in\mathcal{U}}d(\Phi(1,\mathbf{u},H(\theta^{-1}(\mathbf{u}))),\Phi(1,\mathbf{u},F(\theta^{-1}(\mathbf{u}))))$
$=\sup_{\mathbf{u}\in\mathcal{U}}d(g(\mathbf{u}_{1},\Phi(0,\mathbf{u},H(\theta^{-1}(\mathbf{u})))),g(\mathbf{u}_{1},\Phi(0,\mathbf{u},F(\theta^{-1}(\mathbf{u}))))$
$=\sup_{\mathbf{u}\in\mathcal{U}}d(g(\mathbf{u}_{1},H(\theta^{-1}(\mathbf{u}))),g(\mathbf{u}_{1},F(\theta^{-1}(\mathbf{u}))))$
$\leq\sup_{\mathbf{u}\in\mathcal{U}}\mu\cdot d(H(\theta^{-1}(\mathbf{u})),F(\theta^{-1}(\mathbf{u})))$
$=\mu\sup_{\mathbf{u}\in\mathcal{U}}d(H(\theta^{-1}(\mathbf{u})),F(\theta^{-1}(\mathbf{u})))$
$=\mu\rho(H,F).$
So $\mathbf{\Psi}$ is a $\mu$-contraction on $C(\mathcal{U},X)$. Since $C(\mathcal{U},X)$ is a complete metric space, by Banach fixed point theorem, $\mathbf{\Psi}$ has a unique fixed point in $C(\mathcal{U},X)$. That is there exist a unique continuous $H\in C(\mathcal{U},X)$ such that $\mathbf{\Psi}(H)=H$. ∎
Remark 4.
If we can find $\eta>0$ and $10<\mu<$ 1 such that $g$ satisfies
$||g(u,x)-g(v,z)||\leq\eta||u-v||+\mu||x-z||\>\forall\>u,v\in U,\forall x,z\in X.$ (13)
Then one can show that $H$ is continuous with respect to the uniform topology on $\mathcal{U}$ (see Proposition 5.1).
Theorem 3.2.
Under the assumption of Theorem 3.1, the sequence $\mathbf{x}=\{H(\theta^{n-1}(\mathbf{u}))\}_{n\in\mathbb{Z}}$ is the unique entire solutions associated to $\mathbf{u}\in\mathcal{U}$.
Proof.
Let $n,m\in\mathbb{Z},n\geq m$. Then $\mathbf{x}[n]=H(\theta^{n-1}(\mathbf{u}))$. We first show that for any $k\in\mathbb{N}_{0}$,
$\mathbf{x}[n]=\Phi(k,\theta^{n-k}(\mathbf{u}),\mathbf{x}[n-k]).$ (14)
We argue by induction on $k$. The case $k=0$ follows from the definitions of a cocycle. For $k=1$,
$\Phi($ $1,\theta^{n-1}(\mathbf{u}),\mathbf{x}[n-1])$
$=\Phi(1,\theta^{n-1}(\mathbf{u}),H(\theta^{n-2}(\mathbf{u}))),$
$=H(\theta^{n-1}(\mathbf{u}))=\mathbf{x}[n].$
Assume (14) holds $k\geq 0$. Let $\mathbf{z}=\{H(\theta^{n-1}(\mathbf{v})\}_{n\in\mathbb{Z}}$ be the entire solution associated to $\mathbf{v}=\theta^{-1}(\mathbf{u})$. Then
$\Phi(k+1,\>$ $\theta^{n-(k+1)}(\mathbf{u}),\mathbf{x}[n-(k+1)])$
[…]
$=\Phi(1,\theta^{n-1}(\mathbf{u}),\Phi(k,\theta^{n-(k+1)}(\mathbf{u}),H(\theta^{n-1-(k+1)}(\mathbf{u})),$
$=\Phi(1,\theta^{n-1}(\mathbf{u}),\Phi(k,\theta^{n-k}(\mathbf{v}),H(\theta^{n-k-1}(\mathbf{v})))),$
$=\Phi(1,\theta^{n-1}(\mathbf{u}),\Phi(k,\theta^{n-k}(\mathbf{v}),\mathbf{z}[n-k])),$
$=\Phi(1,\theta^{n}(\mathbf{v}),\mathbf{z}[n])\quad\text{by the inductive hypothesis,}$,
$=\mathbf{z}[n+1],$
$=H(\theta^{n}(\mathbf{v}))=H(\theta^{n-1}(\mathbf{u}))=\mathbf{x}[n].$
Choosing $k=n-m$, we have that
$\mathbf{x}[n]=\Phi(n-m,\theta^{m}(\mathbf{u}),\mathbf{x}[m]).$
Remark 5.
Under the assumption of Theorem 3.1, the $\mathbf{u}$-fiber $A_{\mathbf{u}}$ of the pullback attractor is given by $A_{\mathbf{u}}=\{H\circ\theta^{-1}(\mathbf{u})\}$. Furthermore, the pullback attraction of $\{A_{\mathbf{u}}\}_{\mathbf{u}\in\mathcal{U}}$ is uniform since for any $x_{0}\in X$
$\sup_{\mathbf{u}\in\mathbf{U}}d(\Phi(n,\theta^{-n}(\mathbf{u}),x_{0}),H\circ\theta^{-1}(\mathbf{u}))$ $=d(\Phi(n,\theta^{-n}(\mathbf{u}),x_{0}),\Phi(n,\theta^{-n}(\mathbf{u}),H\circ\theta^{-n}(\mathbf{u}))$
$\leq\mu^{n}||x_{0}-H\circ\theta^{-n}(\mathbf{u})||$
$\leq\mu^{n}\text{diam}(X)\to 0\>\text{ as }n\to\infty.$
Remark 6.
The function $H$ is also known as the functional associated with the system
$\mathbf{x}[n]=g(\mathbf{u}[n],\mathbf{x}[n-1])$ (15)
in [11, 14]. In the case that the input sequence are trajectory of a dynamical system, it is referred to as the generalised synchronisation [23, 15]. In this context, it is the function that maps the time evolution of the dynamical system to the time evolution of (15) and as such, a desirable property of the generalised synchronisation is that it is a diffeomorphism onto its image (an embedding in the sense of differential topology). This ensures that the quantities such as Lyapunov exponents and fractal dimension are preserved. The question of whether or not the generalised synchronisation is an embedding is still open in the general case [15] and settled for $g(u,x)=Wx+W^{\text{in}}\omega(u)$ [13] where $\omega$ is a generic one-dimensional observation function of the dynamical systems, $W$ and $W^{\text{in}}$ are generic choices of connection matrices.
3.2 Fractal Dimension of The Pullback Attractor
In this section, we will provide a bound for the dimension of the image of $H$ under certain assumptions on the skew product flow $(\theta,\Phi)$. We will also consider the set
$\mathcal{H}_{\mathcal{V}}=\bigcup_{\mathbf{u}\in\mathcal{V}}H(\mathbf{u})$
for any subset ${\mathcal{V}}\subseteq\mathcal{U}$. For ease of notation, we will denote $\mathcal{H}_{\mathcal{U}}$ as $\mathcal{H}$.
There are various definition of dimension that generalises the notion of dimension on the Euclidean space to metric and topological spaces ([28] provides an excellent introduction to various definition of dimension and their applications in dynamical systems). Key aspects of these generalisation includes
  1. 1.
    the dimension of a singleton is zero,
  2. 2.
    the dimension of an open subset of an $n$-dimensional Euclidean space is $n$.
We will use the upper box-counting dimension or the upper fractal dimension as it provide an upper bound for the other notions of dimension which are use to established bounds on the embedding property of sets. Let $(Z,d)$ be a compact metric space. For any $\varepsilon>0$, let $N(Z,\varepsilon)$ be the minimum number of balls of radius $\varepsilon$ needed to cover $Z$ with centers in $Z$. Note that $N(Z,\varepsilon)$ is finite for each $\varepsilon>0$ since $Z$ is compact. The upper fractal dimension is defined as
$\text{dim}_{B}Z=\lim\sup_{\varepsilon\to 0}\frac{\log_{2}N(Z,\varepsilon)}{-\log_{2}\varepsilon}.$
For nonautonomous dynamical systems, the theory of dimension attempts to show that, under certain conditions, the attractor of a nonautonomous dynamical system is finite dimensional [5, 27, 28, 2]. Dynamics on the attractor or its projection onto the Euclidean space is known as its finite dimensional reduction [7]. For our purposes, we will use the result of Vishik and Chepyzhov to establish upper bounds for the box-counting dimension of $\mathcal{H}_{\mathcal{V}}$ [5, Chapter IX, Theorem 1.2]. We first outline the main assumptions of this result.
For $\mathcal{V}\subseteq\mathcal{U}$, we assume that the skew product flow $(\theta,\Phi)$ satisfies the following assumptions.
Assumption 3.3.
Uniform quasidifferentiability on $\mathcal{V}$: a skew product flow $(\theta,\Phi)$ is uniformly quasidifferentiable on $\mathcal{V}$ if there exist a parameterised family of bounded linear operators $\{L\Phi_{x}(n,\mathbf{u},x)\}_{\mathbf{u}\in\mathcal{V},n\in\mathbb{N}_{0}}$ from $\mathbb{R}^{N_{r}}$ to $\mathbb{R}^{N_{r}}$ such that
$||\Phi(n,\mathbf{u},x)-\Phi(n,\mathbf{u},z)-L_{x}\Phi(n,\mathbf{u},x)(x-z)||\leq\gamma(n,x-z)||x-z||$
where $\mathbf{u}\in\mathcal{V}$, $x,z\in X$ and $\gamma(n,s)\to 0$ as $||s||\to 0$ for all $n\in\mathbb{N}\cup\{0\}$. The function $\gamma(n,s)$ is independent of $x,z\in X$ and $\mathbf{u}\in\mathcal{V}$.
Assumption 3.4.
finite time Lipschitz continuous on $\mathcal{V}$ $i.e$ there exist a function $C:\mathbb{N}\to\mathbb{R}_{+}$ and $v_{0}\in U$ fixed such that for all $\mathbf{u},\mathbf{v}\in\mathcal{V}$ and $x\in X$
$||\Phi(n,\mathbf{u},x)-\Phi(n,\mathbf{v},x)||\leq C(n)||\mathbf{u}_{[0,n]}-\mathbf{v}_{[0,n]}||$
where $\mathbf{u}_{[0,n]}[i]=\mathbf{u}[i]$ for $0\leq i\leq n$ and $\mathbf{u}_{[0,n]}[i]=v_{0}$ otherwise.
Under these assumptions, the following is a simple corollary of [5, Chapter IX, Theorem 1.2] on the box-counting dimension of $\mathcal{H}_{\mathcal{V}}$.
Theorem 3.5 (Vishik and Chepyzhov [5]).
Suppose the conditions of theorem 3.2 holds and that the skew product flow satisfies assumptions 3.3 and 3.3, then the box-counting dimension of $\mathcal{H}_{\mathcal{V}}$ satisfies
[…]
where the box-counting dimension of $\mathcal{V}$ is taken with respect to the product topology.
We will now show that assumption 3.3 and 3.4 are satisfied by the skew product flow $(\theta,\Phi)$ defined in 2.1 if $g$ is twice uniformly differentiable and condition (13) is satisfied.
Lemma 3.6.
Suppose that $X$ is compact and the closure of an open set, $g=(g^{1},\ldots,g^{N_{r}})$ is twice continuously differentiable on $U\times X$ and satisfies condition (13). Then skew product flow $(\theta,\Phi)$ satisfies assumption 3.3 and assumption 3.4.
Proof.
Let $u\in U$ be fixed, and $x,z\in\text{int}X$ be two distinct points. By Taylor’s theorem, there exist $y\in X$ such that
$g^{k}(u,x)-g^{k}(u,z)=\sum_{i=1}^{N_{r}}g^{k}_{x_{i}}(u,x)(x_{i}-z_{i})+\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{r}}g^{k}_{x_{i}x_{j}}(u,y)(x_{i}-z_{i})(x_{j}-z_{j})$
where
$g_{x_{i}}^{k}(u,x)=\frac{\partial g^{k}(u,x)}{\partial x_{i}}\>\>\text{and }\>g_{x_{i}}^{k}(u,x)=\frac{\partial g^{k}(u,x)}{\partial x_{i}\partial x_{j}}.$
Hence
$\frac{||g(u,x)-g(u,z)-D_{x}g(u,x)(x-z)||}{||x-z||}$ $\leq\frac{1}{||x-z||}\max_{k=1,\ldots,N_{r}}\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{r}}|g^{k}_{x_{i}x_{j}}(u,y)(x_{i}-z_{i})(x_{j}-z_{j})|$
$\leq C(u,y)\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{r}}\frac{|x_{i}-z_{i}|\cdot|x_{j}-z_{j}|}{||x-z||}$
$\leq C(u,y)N_{r}\sum_{i=1}^{N_{r}}|x_{i}-z_{i}|\>\>\text{since }\frac{|x_{j}-z_{j}|}{||x-z||}\leq 1$
where $C(u,y)=\max_{k,j,i=1,\ldots,N_{r}}\sum_{i=1}^{N_{r}}\sum_{j=1}^{N_{r}}|g^{k}_{x_{i}x_{j}}(u,y)|$ is a function from $U\times\text{int}X$ to $\mathbb{R}$. Since $g$ is twice continuously differentiable on $U\times X$ and $U\times X$ is compact in $\mathbb{R}^{N_{\text{in}}+N_{r}}$, $C(u,y)$ is bounded with $C=\sup_{(u,y)\in U\times X}C(u,y)$. We take $\gamma(n,\xi)=C\cdot N_{r}\sum_{i=1}^{N_{r}}|\xi_{i}|$.
For assumption 3.4, we note that by cocycle property 10 and assumption (13)
$||\Phi(n,\mathbf{u},x)-\Phi(n,\mathbf{v},x)||$ $\leq\eta||\mathbf{u}[n]-\mathbf{v}[n]||+\mu||\Phi(n-1,\mathbf{u},x)-\Phi(n-1,\mathbf{v},x)||$
$\leq\eta||\mathbf{u}[n]-\mathbf{v}[n]||+\mu\eta||\mathbf{u}[n-1]-\mathbf{v}[n-1]||+\mu^{2}||\Phi(n-2,\mathbf{u},x)-\Phi(n-2,\mathbf{v},x)||$
[…]
$\leq\eta\sum_{i=1}^{n}\mu^{i-1}||\mathbf{u}[n-i]-\mathbf{v}[n-i]||$
$\leq\eta\left(\sum_{i=0}^{n}\mu^{i-1}2^{n-i}\right)\max_{j=0,\ldots n}2^{-j}||\mathbf{u}[j]-\mathbf{v}[j]||$
$=C(n)||\mathbf{u}_{[0,n]}-\mathbf{v}_{[0,n]}||.$
Remark 7.
Theorem 3.5 and Lemma 3.6 implies that
  1. 1.
    $\text{dim}_{B}\>\mathcal{H}_{\mathcal{V}}\leq N_{\text{in}}$ if $\>\mathcal{V}$ is the set of trajectories of an invertible dynamical system on an $N_{\text{in}}$ dimensional Euclidean space (Proposition 5.2),
  2. 2.
    if $M$ is a compact $m$-dimensional manifold, $\omega:M\to\mathbb{R}$ is a generic continuous function, $\phi:M\to M$ is a diffeomorphism and $\mathcal{V}$ is the set all of sequences of the form
    $(\ldots,\omega(\phi^{-2}(u)),\omega(\phi^{-1}(u)),\omega(u),\omega(\phi(u)),\omega(\phi^{2}(u)),\ldots),\>u\in M;$
    then $\text{dim}_{B}\>\mathcal{H}_{\mathcal{V}}\leq m$ (Proposition (5.3).
In the general case, $\text{dim}_{B}\>\mathcal{U}=\infty$ where $\mathcal{U}=U^{\mathbb{Z}}$ for some open subset $U$ of $\mathbb{R}^{N_{\text{in}}}$. As $\mathcal{H}_{\mathcal{U}}\subseteq\mathbb{R}^{N_{r}}$, we already know that $\text{dim}_{B}\>\mathcal{H}_{\mathcal{U}}\leq N_{r}$. In Theorem 3.8, we use the implicit functions theorem [1, Theorem 17.3] to show that $\mathcal{H}$ contains an open set and thus has dimension $N_{r}$. Hence however large one chooses $N_{r}$, the box-counting dimension of $\mathcal{H}$ will always be $N_{r}$. This result is perhaps not surprising since $H(\mathbf{u})$ depends on infinite past histories of $\mathbf{u}$ from the first term where $\mathbf{u}$ takes all possible values in an infinite dimensional metric space $\mathcal{U}=U^{\mathbb{Z}}$. Additionally, in proofs of universal approximations of echo state networks (ESNs) in the class of fading memory functionals, outline in [11], one tries to approximate to uniformly approximate input sequences by a finite dimensional projection using a neural network. The dimensionality of this projection is control by the number of output neurons of the neural network. The required functional on $\mathcal{U}$ is then approximated on the image of this finite dimensional projection. The dimensionality of this the image of this projection increases as the precision of the approximation increases.
Theorem 3.7 (Implicit functions theorem).
Let $V\subseteq\mathbb{R}^{n}$ and $X\subseteq\mathbb{R}^{m}$ be open sets and $G:V\times X\to\mathbb{R}^{m}$ is continuously differentiable function on $V\times X\subseteq\mathbb{R}^{n}\times\mathbb{R}^{m}$. Suppose $G(u_{0},x_{0})=0$ for some $u_{0}\in V$ and $x_{0}\in X$, and the Jacobian $D_{x}G(u_{0},x_{0})$ is invertible. Then there exist an open set $W\subseteq\mathbb{R}^{n}$ containing $u_{0}$ and a unique continuously differentiable function $h:W\to\mathbb{R}^{m}$ such that
  1. 1.
    $h(u_{0})=x_{0}$ and $G(u,h(u))=0$ for every $u\in W$,
  2. 2.
    The Jacobian of $h$ is given by
    $D_{u}h(u)=-[D_{x}G(u,h(u))]^{-1}[D_{u}G(u,h(u))].$
Now we state the main theorem of this paper.
Theorem 3.8.
Let $V\subseteq U$ and $Z\subseteq X$ be open sets. Suppose $g:U\times X\to X$ is uniformly a $\mu$ contraction in the second argument $i.e$ the state variable. Define $\mathcal{U}=U^{\mathbb{Z}}$. Let $m\in\mathbb{N}$ be such that $mN_{\text{in}}\geq N_{r}$. Let $u=(u_{1},u_{2},\ldots,u_{m})\in U^{m}\subseteq\mathbb{R}^{m\times N_{\text{in}}}$ and the function $F:U^{m}\times X\to\mathbb{R}^{N_{r}}$ defined by
$F(u,x)=g(u_{1},g(u_{2},\ldots,g(u_{m},x)\ldots)),\>u=(u_{1},u_{2},\ldots,u_{m})\in\mathbb{R}^{m\times N_{\text{in}}}$ (16)
satisfies the following conditions
  1. 1.
    $F$ is continuously differentiable on $U^{m}\times X$,
  2. 2.
    there exist $z\in Z,v\in V^{m}$ such that $z=F(v,z)$ and the linear map $D_{u}F(v,z):\mathbb{R}^{m\times N_{\text{in}}}\to\mathbb{R}^{N_{r}}$ is surjective $i.e$ it has rank $N_{r}$ and
  3. 3.
    $||D_{x}g(u_{j},x)||_{2}<1$ for all $(u_{j},x)\in U\times X$ and for all $j=1,\ldots,m$.
Then
$\text{dim}_{B}\mathcal{H}=N_{r}.$
The following lemmas will be used in the proof. Lemma 3.9 is an immediate consequence of the submersion theorem [6, Corollary 2.4.9] and it states that submersions are open mappings. See also [29, Theorem 2.1] of the same result under weaker assumptions.
Lemma 3.9.
Let $m\geq n$, $V$ be an open subset of $\mathbb{R}^{m}$ and $h:V\to\mathbb{R}^{n}$ be a continuously differentiable function. Assume that $D_{u}h(u)$ is surjective $i.e$ it has rank $n$ for all $u\in V$, then $h(V)$ is open in $\mathbb{R}^{n}$.
Proof.
Let $u\in V$. By the submersion theorem, there is an open neighbourhood $\widetilde{V}\subseteq V$ of $u$ in $\mathbb{R}^{m}$ and open neighbourhood $\widetilde{W}\supseteq h(\widetilde{V})$ of $h(u)$ in $\mathbb{R}^{n}$ and diffeomorphisms $\phi:\widetilde{V}\to\mathbb{R}^{m}$ and $\varphi:\widetilde{W}\to\mathbb{R}^{n}$ such that the smooth change of variable $\tilde{h}=\varphi\circ h\circ\phi^{-1}:\phi(\widetilde{V})\to\mathbb{R}^{n}$ is a projection onto the first $n$ coordinates. That is, for $x=(x_{1},\ldots,x_{n},\ldots,x_{m})\in\phi(\widetilde{V})$, we have that
$\tilde{h}(x)=h(x_{1},\ldots,x_{n},\ldots,x_{m})=(x_{1},\ldots,x_{n})$
Hence $\tilde{h}\circ\phi(\widetilde{V})$ is open in $\mathbb{R}^{n}$ since $\phi(\widetilde{V})$ is open in $\mathbb{R}^{m}$. Consequently $h(\widetilde{V})=\varphi^{-1}\circ\tilde{h}\circ\phi(\widetilde{V})$ is open since $\varphi$ is a diffeomorphism. So $h(\widetilde{V})$ is an open set in $\mathbb{R}^{n}$ contained in $h(V)$ and containing $h(u)$. ∎
Proof of Theorem 3.8.
We first note that by the assumption on $g$, the map $F$ is uniformly a $\mu^{m}$ contraction in $x$ and thus for any fixed $w\in U^{m}$, there exist a unique $y\in\mathbb{R}^{N_{r}}$ such that $F(w,y)=y$. From Theorem 3.2, $y=H(\theta^{km}(\mathbf{w}))$ for each $k\in\mathbb{Z}$ where $\mathbf{w}$ is the periodic input sequence
$(\ldots,w_{m},w_{m-1},\ldots,w_{1},w_{m},w_{m-1},\ldots,w_{1},\ldots).$
So for any subset $W\subseteq U^{m}$, the set $\mathcal{H}_{\mathcal{W}}$, $\mathcal{W}=W^{\mathbb{Z}}$ of all such $y$’s as $w$ takes values in $W$ is contained in $\mathcal{H}$. In particular, $\text{dim}_{B}\>\mathcal{H}\geq\text{dim}_{B}\>\mathcal{H}_{\mathcal{W}}$.
Now let $G(u,x)=x-F(u,x)$. Then $G$ is continuously differentiable on $U^{m}\times X$ with
$D_{x}G(u,x)=\mathbb{I}-D_{x}g(u_{1},y_{1})D_{x}g(u_{2},y_{2})\ldots D_{x}g(u_{m-1},y_{m-1})D_{x}g(u_{m},y_{m})$
where
$y_{j}=g(u_{j+1},y_{j+1})\>\text{ for }j<m\>\text{ and }\>y_{m}=x.$ (17)
Since
$||D_{x}g(u_{1},y_{1})$ $D_{x}g(u_{2},y_{2})\ldots D_{x}g(u_{m-1},y_{m-1})D_{x}g(u_{m},y_{m})||_{2}$
$\leq||D_{x}g(u_{1},y_{1})||_{2}\cdot||D_{x}g(u_{2},y_{2})||_{2}\cdot\ldots\cdot||D_{x}g(u_{m-1},y_{m-1})||_{2}\cdot||D_{x}g(u_{m},y_{m})||_{2}$
$<1,$
$D_{x}G(u,x)$ is invertible for all $x\in X,u\in U^{m}$ with
$[D_{x}G(u,x)]^{-1}=\sum_{k=0}^{\infty}[D_{x}g(u_{1},y_{1})$ $D_{x}g(u_{2},y_{2})\ldots D_{x}g(u_{m-1},y_{m-1})D_{x}g(u_{m},y_{m})]^{k}.$
Since $G(v,z)=0$, by the implicit functions theorem, there exists an open neighbourhood $V_{0}$ of $v$ and a unique continuously differentiable function $h:V_{0}\to X$ such that $h(v)=z$ and
$G(u,h(u))=0$
for all $u\in V_{0}$. We first note that $h(V_{0})\subseteq\mathcal{H}$. Since $D_{u}F(z,v)$ has rank $N_{r}$, there is an $N_{r}\times N_{r}$ submatrix of $D_{u}F(v,z)$ which is invertible. Thus there exist a neighbourhood $X_{1}$ of $z$ and $V_{1}$ of $v$ such that for all $x\in X_{1}$ and $u\in V_{1}$, the corresponding submatrix of $D_{u}F(u,x)$ remains invertible. So $D_{u}F(u,x)$ has rank $N_{r}$ for all $x\in X_{1}$ and $u\in V_{1}$. Take $W=h^{-1}(X_{1})\cap V_{0}\cap V_{1}$ which is an open neighbourhood of $v$ satisfying $h(W)\subseteq X_{1}$. Hence $D_{u}F(h(u),u)$ has rank $N_{r}$ and $G(u,h(u))=0$ for all $u\in W$. Thus the linear map
$D_{x}h(u)=-[D_{x}G(u,h(u))]^{-1}[D_{u}G(u,h(u))]=-[D_{x}G(u,h(u))]^{-1}[D_{u}F(u,h(u))],$
has rank $N_{r}$ on $W$. Consequently, by Lemma 3.9, $h(W)\subseteq\mathcal{H}_{\mathcal{W}}$ is open in $\mathbb{R}^{N_{r}}$ and thus has box-counting dimension $N_{r}$. Thus
$\text{dim}_{B}\>\mathcal{H}\geq\text{dim}_{B}\>h(W)=N_{r}.$
The space of all input sequences of the type $\mathcal{W}=W^{\mathbb{Z}}$ where $W\subseteq\mathbb{R}^{m\times N_{\text{in}}}$ for all $m\in\mathbb{N}$ can in fact be used to approximate any point in $\mathcal{H}$.
Theorem 3.10.
The set
$\mathcal{H}^{N_{r}}:=\{z\in X:\exists\>m\in\mathbb{N}\>\text{ and}\>u\in U^{mN_{\text{in}}}\>\text{ with }\>F(z,u)=z\text{ and }\text{rank}\>D_{u}F(z,u)=N_{r}\}$
is dense in $\mathcal{H}$.
The following subset of input sequences will be used in the proof.
Definition 3.1.
For any $m\in\mathbb{N}$, let
$P(\mathcal{U},m)=\{\mathbf{u}\in\mathcal{U}:\mathbf{u}[km+r]=\mathbf{u}[r]\>\forall\>k\in\mathbb{Z},r=-m+1,\ldots,0\}$
The set $P$ contains all $m$-periodic input sequences whose terms from $-m+1$ to 00 repeats from $1$ to $m$, then from $m+1$ to $2m$ and so on. Similarly for negative indices, the terms from $-m+1$ to 00 repeats from $-2m$ to $-m$ and so on. In the proof, we use the following two facts about $P(\mathcal{U},m)$
  1. 1.
    $P(\mathcal{U},m)$ is invariant under $\theta$ i.e $\theta(P(\mathcal{U},m))=\theta^{-1}(P(\mathcal{U},m))=P(\mathcal{U},m)$.
  2. 2.
    $P(\mathcal{U},m)$ is closed subset of $\mathcal{U}$ and hence compact. This is because a sequence $\{\mathbf{u}_{n}\}_{n=1}^{\infty}$ converges to $\mathbf{u}$ in the product topology if and only if the terms $\mathbf{u}_{n}[k]\to\mathbf{u}[k]$ as $n\to\infty$ for each $k\in\mathbb{Z}$. So a sequence of $m$-periodic input sequences will converge to an $m$-periodic input sequence.
We will use the following fact that singular matrices are isolated.
Lemma 3.11.
Consider the $n\times n$ matrix $A$ as subset of $\mathbb{R}^{n^{2}}$ with entries $a_{ij}\in\mathbb{R}$. Suppose $A$ is singular. For any $\varepsilon>0$, there exist an open set $\mathbf{B}\subseteq\mathbb{R}^{n^{2}}$ such any $n\times n$ matrix $B\in\mathbf{B}$ is invertible and satisfies
$0<\max_{i,j=1\ldots,n}|a_{ij}-b_{ij}|<\varepsilon.$
Proof.
The proof is an immediate consequence of [17, Theorem 1.2.17] which allows us to perturb the diagonal entries of $A$ to get a matrix whose determinant is non-zero and the result follows from the continuity of the determinant. ∎
Proof of Theorem 3.10.
We will prove the following stronger statement:
$\lim_{m\to\infty}\text{dist}(\mathcal{H},H(P(\mathcal{U},m)))=0.$
Since $\mathcal{U}$ is compact, $H$ is uniformly continuous. Let $\varepsilon>0$ and choose $\delta>0$ so that for all $\mathbf{v},\mathbf{u}\in\mathcal{U}$
$D(\mathbf{u},\mathbf{v})\leq\delta\implies||H(\mathbf{u})-H(\mathbf{v})||\leq\frac{\varepsilon}{2}.$
We choose $n$ large enough so that $2^{-n}\leq\delta$. So for $|j|>n$, we have that
$2^{-|j|}<2^{-n}\leq\delta.$
Let $m=2n+1$. Now for any $\mathbf{u}\in\mathcal{U}$, let $\mathbf{v}\in P(\mathcal{U},m)$ be the input sequence
$\mathbf{v}=(\ldots,\mathbf{u}[0],\mathbf{u}[-2n],\ldots,\mathbf{u}[-n],\ldots,\mathbf{u}[-1],\mathbf{u}[0],\mathbf{u}[-2n],\ldots)$
where the terms of $\mathbf{v}$ satisfies $\mathbf{v}[j]=\mathbf{u}[r]$ where $j=r+mk$ for some $k\in\mathbb{Z}$ and $r=-2n,-2n+1,\ldots,0$. The input sequences $\theta^{-n}(\mathbf{u})$ and $\theta^{-n}(\mathbf{v})$ satisfies
$D(\theta^{-n}(\mathbf{u}),\theta^{-n}(\mathbf{v}))$ $=\sup_{j\in\mathbb{Z}}2^{-|j|}d(\mathbf{u}[j-n],\mathbf{v}[j-n])$
$\leq\max\{2^{-n},2^{-n}d(\mathbf{u}[-2n],\mathbf{v}[-2n]),\ldots,d(\mathbf{u}[-n],\mathbf{v}[-n]),\ldots,2^{-n}d(\mathbf{u}_{0},\mathbf{v}_{0})\}$
$=2^{-n}\leq\delta.$
Hence
$||H(\theta^{-n}(\mathbf{u}))-H(\theta^{-n}(\mathbf{v}))||\leq\frac{\varepsilon}{2}.$
Since $\theta^{-n}(\mathcal{U})=\mathcal{U}$ and $\theta^{-n}(P(\mathcal{U},m))=P(\mathcal{U},m)$, for any $\mathbf{u}\in\mathcal{U}$, there exist $\mathbf{v}\in P(\mathcal{U},m)$ so that
$||H(\mathbf{u})-H(\theta^{-1}(\mathbf{v}))||\leq\frac{\varepsilon}{2}.$
Note that $H(\theta^{-1}(\mathbf{v}))$ is the fixed point of the map $F:X\times U^{m}\to X$ associated with the input $\mathbf{v}^{m}=(\mathbf{v}[-2n],\ldots,\mathbf{v}[-n],\ldots,\mathbf{v}[0])\in U^{m}$. Suppose
$\text{rank}\>D_{u}F(H(\theta^{-1}(\mathbf{u})),\mathbf{v}^{m})<N_{r}.$
Using Lemma 3.11 and the implicit functions theorem 3.7, let $\mathbf{u}^{m}=(\mathbf{u}^{m}[-2n],\ldots,\mathbf{u}^{m}[-n],\ldots,\mathbf{u}^{m}[0])$ be a point near $\mathbf{v}^{m}$ in $\mathbb{R}^{mN_{\text{in}}}$ and $z$ be near $H(\theta^{-1}(\mathbf{v}))$ such that
$z=F(z,\mathbf{u}^{m}),\quad||H(\theta^{-1}(\mathbf{v}))-z||\leq\frac{\varepsilon}{2}\>\text{ and }\>\text{rank}\>D_{u}F(z,\mathbf{u}^{m})=N_{r},$
where $\bar{\mathbf{u}}$ is $m$-periodic input sequences whose first term is $\mathbf{u}^{m}[-2n]$. Thus
$||H(\theta^{-1}(\mathbf{u}))-z||\leq||H(\theta^{-1}(\mathbf{u}))-H(\theta^{-1}(\mathbf{v}))||+||H(\theta^{-1}(\mathbf{v}))-z||\leq\varepsilon.$
3.2.1 Application to Echo State Networks
If $g$ is the state update equation of an ESN $i.e$ $g(u_{j},x)=\text{tanh}(Wx+W^{\text{in}}u_{j})$ where we assume that the connectivity matrices $W$ and $W^{\text{in}}$ both have full rank and $||W||_{2}<1$, then the assumptions of Theorem 3.8 holds. The connectivity matrices $W$ and $W^{\text{in}}$ are usually chosen randomly and thus the assumption of full rank holds almost surely. Here
$F(u,x)=\text{tanh}(W^{\text{in}}u_{2}+\text{tanh}(W^{\text{in}}u_{1}+\ldots(W^{\text{in}}u_{m}+Wx))\ldots)$
is differentiable since each of the partial derivatives exist and are continuous. In addition
$||D_{x}g(u_{j},x)||_{2}$ $=||D_{x}\text{tanh}(W^{\text{in}}u_{j}+Wx)||_{2}$
$=||\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{j}+Wx))W||_{2}$
$\leq||\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{j}+Wx))||_{2}\cdot||W||_{2}\leq||W||_{2}<1$
for all $(u_{j},x)\in U\times X$. The Jacobian with respect to $u$ is
$D_{u}F(u,x)$ $=\begin{bmatrix}\frac{\partial F_{1}(u,x)}{\partial u_{11}}&\frac{\partial F_{1}(u,x)}{\partial u_{12}}&\frac{\partial F_{1}(u,x)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{21}}&\frac{\partial F_{1}(u,x)}{\partial u_{21}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{2N_{\text{in}}}}&\ldots&\frac{\partial F_{1}(u,x)}{\partial u_{mN_{\text{in}}}}\\ \vdots&\ddots&\ddots&\ldots&\ddots&\ddots&\dots&\ddots&\ldots&\vdots\\ \frac{\partial F_{N_{r}}(u,x)}{\partial u_{11}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{12}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{1N_{\text{in}}}}&\ldots&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{21}}&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{2N_{\text{in}}}}&\ldots&\ldots&\frac{\partial F_{N_{r}}(u,x)}{\partial u_{mN_{\text{in}}}}\end{bmatrix}$
$=\begin{bmatrix}S_{1}W^{\text{in}}&|&S_{2}W^{\text{in}}&|&\ldots&|&S_{m}W^{\text{in}}\end{bmatrix}$
$=\begin{bmatrix}S_{1}&S_{2}&&\ldots&S_{m}\end{bmatrix}\begin{bmatrix}W^{\text{in}}&\mathbf{0}&\ldots&\mathbf{0}\\ \mathbf{0}&W^{\text{in}}&\ldots&\mathbf{0}\\ \vdots&\ldots&\ddots&\vdots\\ \mathbf{0}&\mathbf{0}&\ldots&W^{\text{in}}\end{bmatrix}$
where
$S_{j}=\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{1}+Wy_{1}))\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{2}+Wy_{2}))\ldots\text{diag}(\operatorname{sech}^{2}(W^{\text{in}}u_{j}+Wy_{j}))$
and the $y_{j}$’s are defined in (17). Hence if we choose $m$ so that $mN_{\text{in}}\geq N_{r}$, we have that for all $u\in U^{m}$ and $x\in X$, $D_{u}F(u,x)$ is full rank. Thus by Theorem 3.8
$\text{dim}_{B}\mathcal{H}=N_{r}.$
4 Conclusions
In this paper, we provide two important results for the box-counting dimension of the pullback attractor of reservoir systems. First, using a result of Vishik and Chepyzhov [5], we demonstrate that the box-counting dimension of the pullback attractor of a driven reservoir system is bounded above by the box-counting dimension of the space of input sequences in the product topology. A key ingredient to extending this result to a general embedding theorem for RC systems is to understand the properties of space of trajectories of invertible dynamical systems in the product topology. As the skew product flow is dynamical system of the product space $\mathcal{U}\times X$, a version of Takens’ delay embedding theorem [27] applies to their compact invariant sets with finite box-counting dimension. In particular, if one can establish an embedding between a dynamical system and its space of trajectories in the product topology, one can lift the question of whether or the not the generalised synchronisation is an embedding into the corresponding question for the skew product flow. This approach, if it works, will avoid the intricacies of working for example on a compact manifold.
In the infinite dimensional case, we show that the pullback attractor always contains an open set and thus has dimension equal to that of the driven reservoir. The proof technique relied on a subset of input sequences composed periodic input sequences whose period is larger that dimension of the reservoir. This essentially reduces the problem to a fixed point problem where powerful tools such as the implicit functions theorem and submersion theorem can be applied.
Acknowledgements
The author thanks Peter Ashwin and George Datseris for their insightful comments. This is supported by EPSRC studentship funding via EP/W523859/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
5 Appendix
Proof of Lemma 2.1.
Given $\mathbf{u}\in\mathcal{U}$, let $x_{0}\in X$ and $m\in\mathbb{N}_{0}$ be fixed. We are argue by induction on $n$. The case where $n=0$ follows from the (8). Assume that it holds for $n-1$ for $n\in\mathbb{N}$, then
$\Phi(n,\theta^{m}(\mathbf{u}),\Phi(m,\mathbf{u},x_{0}))$ $=g(\pi_{0}(\theta_{n}(\theta_{m}(\mathbf{u}))),\Phi(n-1,\theta^{m}(\mathbf{u}),\Phi(m,\mathbf{u},x_{0})))$
$=g(\pi_{0}(\theta_{m+n}(\mathbf{u})),\Phi(n-1,\theta^{m}(\mathbf{u}),\Phi(m,\mathbf{u},x_{0})))$
$=g(\pi_{0}(\theta_{m+n}(\mathbf{u})),\Phi(m+n-1,\mathbf{u},x_{0}))$
$=\Phi(m+n,\mathbf{u},x_{0}).$
Proposition 5.1.
If there exist positive constants $\mu,\eta$ with $\mu<1$ such that the function $g$ satisfies
$||g(u,x)-g(v,z)||\leq\eta||u-v||+\mu||x-z||.$ (18)
Then for any $\mathbf{u},\mathbf{v}\in\mathcal{U}$,
$||H(\mathbf{u})-H(\mathbf{v})||\leq C\sup_{j\in\mathbb{Z}}||\mathbf{u}_{j}-\mathbf{v}_{j}||$
where $C$ is a constant independent of $\mathbf{u}$ and $\mathbf{v}$.
Proof.
Let $\mathbf{u},\mathbf{v}\in\mathcal{U}$. By Theorem 3.2, $H(\mathbf{u})=\mathbf{x}_{1}=\Phi(1,\mathbf{u},H\circ\theta^{-1}(\mathbf{u}))=g(\mathbf{u}_{1},H\circ\theta^{-1}(\mathbf{u}))$. Hence for any $k\in\mathbb{N}$
$||H(\mathbf{u})-H(\mathbf{v})||$ $=||g(\mathbf{u}_{1},H\circ\theta^{-1}(\mathbf{u}))-g(\mathbf{v}_{1},H\circ\theta^{-1}(\mathbf{v}))||$
$\leq\eta||\mathbf{u}_{1}-\mathbf{v}_{1}||+\mu||H\circ\theta^{-1}(\mathbf{u})-H\circ\theta^{-1}(\mathbf{v})||$
$\leq\eta||\mathbf{u}_{1}-\mathbf{v}_{1}||+\mu\eta||\mathbf{u}_{0}-\mathbf{v}_{0}||+\mu^{2}||H\circ\theta^{-2}(\mathbf{u})-H\circ\theta^{-2}(\mathbf{v})||$
$\vdots$
$\leq\eta\sum_{i=1}^{k}\mu^{i-1}||\mathbf{u}_{2-i}-\mathbf{v}_{2-i}||+\mu^{k}||H\circ\theta^{-k}(\mathbf{u})-H\circ\theta^{-k}(\mathbf{v})||.$
Letting $k\to\infty$, we have that
$||H(\mathbf{u})-H(\mathbf{v})||\leq\eta\sum_{i=1}^{\infty}\mu^{i-1}||\mathbf{u}_{2-i}-\mathbf{v}_{2-i}||$
since $||H\circ\theta^{-k}(\mathbf{u})-H\circ\theta^{-k}(\mathbf{v})||$ remains bounded and $\mu^{k}\to 0$ as $k\to\infty$. Hence
$||H(\mathbf{u})-H(\mathbf{v})||\leq\eta\sum_{i=1}^{\infty}\mu^{i-1}||\mathbf{u}_{2-i}-\mathbf{v}_{2-i}||=\eta\sum_{j=1}^{-\infty}\mu^{1-j}||\mathbf{u}_{j}-\mathbf{v}_{j}||\leq C\sup_{j\in\mathbb{Z}}||\mathbf{u}_{j}-\mathbf{v}_{j}||$
where $C=\eta\sum_{j=1}^{-\infty}\mu^{1-j}=\frac{\eta}{1-\mu}$. ∎
Proposition 5.2.
Suppose $\mathcal{V}\subseteq U^{\mathbb{Z}}$ is the set of all trajectories of an invertible discrete time dynamical system on $U$ where $U$ is a compact subset of $\mathbb{R}^{N_{\text{in}}}$. Then, in the product topology
$\text{dim}_{B}\mathcal{V}\leq N_{\text{in}}.$
Proof.
Let $\varepsilon>0$. We will show that the number of balls of radius $\varepsilon$ needed to cover $\mathcal{V}$ in the product topology is at most $N(U,\varepsilon)$, the number of balls of $\varepsilon$ needed to cover $U$ in $\mathbb{R}^{N_{\text{in}}}$. Let $\{B_{\varepsilon}(u_{i})\}_{i=1}^{N(U,\varepsilon)}$ be a cover of $U$ with $N(U,\varepsilon)$ balls of radius $\varepsilon$ centered at $u_{i}\in U$. By definition of $\mathcal{V}$ there exist exactly one input sequence with whose zeroth term in $u_{i}$. Let $\mathbf{u}^{i}$ be the input sequence passing through $u_{i}$ with $\mathbf{u}^{i}_{0}={u}_{i}$. Now for any $\mathbf{v}\in\mathcal{V}$, we have that
$D(\mathbf{u}^{i},\mathbf{v})=\sup_{j\in\mathbb{Z}}2^{-|j|}d(\mathbf{u}^{i}_{j},\mathbf{v}_{j}))\leq d(\mathbf{u}^{i}_{0},\mathbf{v}_{0})\leq\varepsilon$
for some $i\in 1,\ldots,N(U,\varepsilon)$. Thus $N(\mathcal{V},\varepsilon)\leq N(U,\varepsilon)$ and the result follows from this. ∎
Proposition 5.3.
Let $M$ be a compact manifold of dimension $m$ and let $\phi:M\to M$ be a smooth diffeomorphism and $\omega:M\to\mathbb{R}$ a smooth function. The space $\mathcal{V}$ of all $1$-dimension input sequences of the form
$(\ldots,\omega(\phi^{-2}(u)),\omega(\phi^{-1}(u)),\omega(u),\omega(\phi(u)),\omega(\phi^{2}(u)),\ldots)$
where $u\in M$ satisfies
$\dim_{F}\mathcal{V}\leq m$
in the product topology on $\mathbb{R}^{\mathbb{Z}}$.
Proof.
For $u\in M$, consider the $(2m+1)$-delay map
$\Phi_{(\phi,\omega)}(u)=(\omega(u),\omega(\phi(u)),\ldots,\omega(\phi^{2m}(u))).$
By Taken’s delay embeddings theorem [32], $\Phi_{(\phi,\omega)}$ is an embedding and so $\dim_{F}\Omega=m$ where $\Omega=\Phi_{(\phi,\omega)}(M)$. Let $\Omega$ be considered as subset of $\mathbb{R}^{2m+1}$ endowed with metric
$d_{\infty}(u,v)=\max_{1\leq i\leq 2m+1}|x_{i}-y_{i}|,\quad x,y\in\mathbb{R}^{2m+1}.$
For any $\varepsilon>0$, we can cover $\Omega$ with $N(\varepsilon,\Omega)$ balls of radius $\varepsilon$ with centres $\Phi_{(\phi,\omega)}(u^{i})\in\Omega$, $u^{i}\in M$ and $i=1,\ldots,N(\varepsilon,\Omega)$. As with the proof of Proposition (5.2), we can define the $(2m+1)$-periodic input sequence $\mathbf{u}^{i}$ whose terms from 00 to $2m$ are $\Phi_{(\phi,\omega)}(u^{i})$. From this we have that $N(\varepsilon,\mathcal{V})\leq N(\varepsilon,\Omega)$. ∎